OpenStudy (anonymous):

Determine the value(s) for which the rational expression is undefined. -2y/7y-2 please help someone?

OpenStudy (anonymous):

Here, the denominator cannot be zero.

OpenStudy (anonymous):

The expression is undefined where: 7y - 2 = 0

OpenStudy (anonymous):

what happened to the -2y?

OpenStudy (anonymous):

The -2y does not matter at all. That term can be anything and it would still be defined. Division by zero, however, is not defined.

OpenStudy (anonymous):

y=2 for 7y-2=0?

OpenStudy (anonymous):

No..simply solve 7y-2 = 0 for y. That will be where the function is not defined.

OpenStudy (anonymous):

y=2/7?

OpenStudy (anonymous):

y=-2/7?

OpenStudy (anonymous):

No negative, just y = 2/7. So the function is undefined when y = 2/7

OpenStudy (anonymous):

y=2/7 is not one of my options

OpenStudy (anonymous):

Is it: